# Retro-Expert: Collaborative Reasoning for Interpretable Retrosynthesis

### v1

@@ Abstract
Retrosynthesis prediction aims to infer the reactant molecule based on a given product molecule, which is a fundamental task in chemical synthesis. However, existing models rely on static pattern-matching paradigm, which limits their ability to perform effective logic decision-making, leading to black-box decision-making. Building on this, we propose Retro-Expert, an interpretable retrosynthesis framework that performs collaborative reasoning by combining the complementary reasoning strengths of Large Language Models and specialized models via reinforcement learning. It outputs natural language explanations grounded in chemical logic through three components: (1) specialized models perform shallow reasoning to construct high-quality chemical decision space, (2) LLM-driven critical reasoning to generate predictions and corresponding interpretable reasoning path, and (3) reinforcement learning optimizing interpretable decision policy. Experiments show that Retro-Expert not only surpasses both LLM-based and specialized models across different metrics but also provides expert-aligned explanations that bridge the gap between AI predictions and actionable chemical insights.
Introduction
@@
Retrosynthesis prediction aims to deduce potential reactants and reaction pathways for synthesizing a target product molecule based on its structural characteristics (Somnath et al. 2021; Segler and Waller 2017; Sun et al. 2021), holding significant application value in drug discovery and molecular design (Hu et al. 2025; Wang et al. 2023, 2018). However, existing models primarily adopt a pattern-matching paradigm, which learn mappings between product SMILES and reactant SMILES from datasets, framing the task as either classification or auto-regressive sequence generation (Yao et al. 2023; Chen and Jung 2021; Yao et al. 2023; Zheng et al. 2019). Under this paradigm, traditional specialized models capture intricate structural features to learn improved mappings, yet their end-to-end nature inherently limits their ability to provide chemically-grounded rationales for their predictions. In contrast, large language models (LLMs) rely on supervised fine-tuning (SFT) to enhance retrosynthetic capabilities, but their performance depends heavily on memorized reaction patterns and typically outputs only reactants or corresponding options (Zhang et al. 2024; Zhao et al. 2024). This reliance limits their capacity to understand underlying chemical principles and perform logical reasoning. This paradigm faces three key limitations: (1) The model only generates reactant SMILES strings, with no transparency in its internal reasoning process. (2) The predictions lack natural language explanations grounded in chemical logic, critically hindering real-world adoption. (3) The pure pattern-matching mechanism significantly constrains the model’s capacity for effective logical decision-making. These limitations lead to a lack of reliable basis for the predictions in chemical principles, which severely undermines chemists’ trust in practical applications.
Figure 1: Retro-Expert: an interpretable retrosynthesis framework capable of generating step-by-step reasoning in natural language, enabled by the collaboration the reasoning capabilities of LLMs with the knowledge grounding capabilities of specialized models.
@@
Given the mentioned challenges, a reasoning-driven paradigm aligns more closely with the intrinsic nature of retrosynthetic analysis than simplistic pattern matching. This approach mirrors the strategy employed by chemical experts, who conceptualize retrosynthesis as an iterative and logical reasoning process. Notably, recent breakthroughs in other domains (Wang et al. 2025; Xie et al. 2025; Chen et al. 2025) have demonstrated the potential of LLMs to address complex specialized problems through specialized-knowledge-based reasoning, enhanced by reinforcement learning. These findings motivate us to explore how to leverage LLMs’ emergent reasoning capabilities to enhance the interpretability of retrosynthesis prediction.
@@
We focus on chemical knowledge-based retrosynthetic reasoning by LLMs to generate reactants along with explainable reasoning process, ensuring interpretable and transparent retrosynthesis prediction. To induce reasoning, the selection of an appropriate learning paradigm is crucial. Supervised Fine-Tuning (SFT) inclines models towards replicating prevalent reaction patterns observed in training data, rather than engaging in reasoning grounded in underlying chemical principles. In contrast, Reinforcement Learning (RL) approaches, which optimize models’ behavioral policy through reward feedback, offer a theoretically viable path for models to learn reasoning within complex decision spaces.
@@
However, directly applying RL to incentivize models for retrosynthetic reasoning faces two critical challenges: (1) Domain Knowledge Disparity. Retrosynthesis demands not only logical reasoning but also mastery of specialized chemical knowledge. Pre-trained LLMs fail to adequately internalize and apply specific chemical principles when reasoning solely based on molecular SMILES. (2) Granularity Mismatch in Specialized Capabilities. The retrosynthetic reasoning process comprises multiple subtasks with distinct cognitive requirements, including reaction type classification and reaction center localization (Gao et al. 2022; Wang et al. 2021; Yan et al. 2020). While general LLMs demonstrate profound logical reasoning abilities, they inherently lack precise perception and matching capabilities for fundamental chemical reaction patterns - a critical requirement for certain subtasks. In contrast, dedicated specialized models, each individually optimized for a specific subtask, have achieved expert-level performance (Wang et al. 2023; Chen and Jung 2021; Somnath et al. 2021). To bridge this gap, we propose a novel paradigm that synergistically combines: (1) the experience-based shallow reasoning (pattern recognition) of specialized models with (2) the deep logical reasoning of LLMs. Specifically, specialized models construct high-quality, multi-dimensional chemical decision spaces, upon which LLMs perform interpretable logical reasoning.
@@
Building upon these insights, we present Explainable and Cooperative retrosynthesis framework, Retro-Expert, the first explainable retrosynthesis framework that integrates natural language-based expert reasoning with model-agnostic compatibility. The framework contains three synergistic core components: (1) Chemical Decision Space Construction. Leveraging specialized models (e.g., reaction type classifiers, reaction center localization and reactant generators) to construct high-quality chemical decision space, which provides “knowledge anchors” for LLM’s subsequent deep reasoning. (2) Collaborative Reasoning Engine. Leveraging the effective collaboration between shallow pattern recognition and deep logical reasoning, the LLM perform integrated critic analysis and in-depth reasoning upon the chemical decision space. It produces the final reactant prediction along with interpretable, natural language-based reasoning process. (3) Knowledge-Guided Policy Optimization. Retro-Expert optimizes the LLM’s reasoning strategy via rule-based reinforcement learning. A multi-stage reward mechanism is designed during training to guide the model toward learning an optimal and trustworthy reasoning path. Notably, Retro-Expert allows seamless integration of arbitrary specialized models during inference, enabling flexible expansion without retraining.
@@
2. We propose a collaborative retrosynthesis reasoning framework built upon the synergy between the LLM and specialized models. It improves prediction accuracy while generates human-understandable step-by-step analysis.
@@
Related Work
Retrosynthesis Prediction. Existing retrosynthesis methods are broadly categorized into two modeling paradigms: specialized machine learning (ML) models and large language model (LLM)-based methods. ML-based approaches model chemical molecules from different perspectives (e.g., SMILES strings, molecular graphs) to learn the mapping from product to reactant molecules (Sacha et al. 2021; Zhong et al. 2022; Chen and Jung 2021; Somnath et al. 2021). Although these methods demonstrate significant potential through precise molecular mappings, their black-box decision-making mechanisms inherently restrict the model’s interpretability.
LLM-based approaches aim to enhance the comprehension and memorization capabilities of LLMs for chemical tasks. Some efforts tailor LLMs to retrosynthesis via SFT (Zhang et al. 2024; Zhao et al. 2024; Yang et al. 2024). Limited by the memorization-centric nature of SFT, these models tend to replicating memorized reaction patterns and only produce predicted reactants or corresponding options. Other work integrates the LLM with external tools to improve prediction accuracy (Bran et al. 2023, 2025; Liu et al. 2024). However, they are limited to directly adopting or scoring based on predefined metrics of specialized model predictions. Such strategies lack engagement in the predictive process and thus fail to provide interpretable reasoning pathways that reveal underlying reaction mechanisms.
@@
Large Language Model Reasoning. In recent years, the profound reasoning capabilities of large language models (LLMs) have been developed to solve specialized scientific problems (Su et al. 2025; Tang et al. 2025; Putri, Taufiq et al. 2025; Pan et al. 2025). While employing SFT (Zhou et al. 2023; Huang et al. 2025), the model primarily learns to replicate common reasoning patterns from curated datasets, restricting its effectiveness on more complex tasks. In contrast, reinforcement learning (RL)-based reasoning models recently achieve significant progress (Guo et al. 2025; Zhou et al. 2024; Feng et al. 2025). In the chemistry field, some studies have assessed the performance of LLMs on chemical reasoning tasks (Guo et al. 2023; Feng et al. 2024). These models predominantly focus on enhancing performance through SFT paradigm. However, there are few studies that incorporate expert-level reasoning strategies to tackle complex chemical reasoning tasks.
Methodology
Overview
Retro-Expert is to enhance prediction accuracy while generating human-understandable reasoning process. As illustrated in Figure 2, Retro-Expert is built upon three core modules. The process starts with Chemical Decision Space Construction, where specialized models perform experience-based shallow reasoning on a target product to construct an anchored chemical decision space. This foundation activates the Collaborative Reasoning Engine (CRE), where the LLM performs deep reasoning by critically analyzing candidates, selecting an optimal path, and articulating its multi-step logic into a complete natural language chain. Finally, Knowledge-Constrained Decision Policy Optimization employs RL to refine the LLM’s decision-making strategy, using a multi-stage reward mechanism to guide it towards an optimal and trustworthy reasoning path.
Problem Formulation
The objective of the retrosynthesis task $T_{\text{retro}}$ is to predict the set of reactants $\{M_{r}^{i}\}_{i=1}^{C}(C\geq 1)$ corresponding to a target product $M_{p}$. According to chemical expert knowledge, the overall task can be decomposed into $n$ logically connected subtasks $T_{\text{retro}}=\{T_{0},T_{1},\dots,T_{n}\}$. Given an environment $\mathcal{E}$ containing $N(N\geq n)$ specialized models $\mathcal{M}=\{m_{0},m_{1},\dots,m_{N}\}$, each model $m_{i}$ performs experience-based domain-specific reasoning to generate Top-$K$ candidate predictions $P_{i}$ for its corresponding subtask:
$P_{i}=\{P_{i}^{k}\}_{k=1}^{K},\quad P_{i}^{k}\sim p(m_{i}|M_{p};\theta),$ (1)
where $\theta$ denotes the parameters of model $m_{i}$, and $P_{i}^{k}$ represents the $k$-th candidate result. The chemical decision space $\mathcal{T}$ is defined as the Cartesian product of candidate results from all subtasks:
$\mathcal{T}=(P_{0},P_{1},\dots,P_{n}),$ (2)
with a space size of $K^{n}$. Upon this space, an LLM $M_{\text{LLM}}$ interacts with $\mathcal{E}$, critically analyzes the Top-$K$ candidates from models in $\mathcal{M}$, and sequentially infers the correct answer $P_{i}^{\prime}$ for each subtask. This generates a reasoning path $\mathcal{T}_{\text{LLM}}=(P_{0}^{\prime},P_{1}^{\prime},\dots,P_{n}^{\prime})$ alongside a natural-language explanation $R$. Retro-Expert has two core objectives: generating the correct set of reactants $\hat{a}=\{M_{r}^{i}\}_{i=1}^{C}$, and identifying the optimal path $T^{*}$ with the highest reward within $\mathcal{T}$:
@@
Chemical Decision Space Construction Module
Retro-Expert introduces a paradigm shift in retrosynthesis by reformulating it from a mapping problem into a multi-stage reasoning task, where we use specialized models to construct a well-defined structured chemical decision space.
@@
Specifically, for a given target product $M_{p}$, we first decompose the retrosynthesis task into a set of distinct chemical sub-tasks such as reaction type prediction and reaction center localization. We then invoke a suite of specialized models, each dedicated to a sub-task, to generate a corresponding set of plausible candidates. This process transforms the implicit heuristics embedded within the specialized models into explicit, actionable chemical information. These candidate sets are then integrated, with each dimension of chemical knowledge (e.g., reaction type, reaction center) forming a multi-dimensional decision space. This space provides well-founded “knowledge anchors” that ground the LLM’s subsequent reasoning, enabling it to progress from heuristic selection to sophisticated, knowledge-aware strategic planning.
The design of this module is guided by two core principles to ensure its effectiveness. First, it prioritizes comprehensive candidate coverage by employing specialized models optimized for high recall. The resulting breadth of options creates a chemically grounded and reliable search space for the LLM, significantly reducing the risk of factual hallucination. Second, the architecture is engineered for modularity and forward compatibility. Being agnostic to the type or number of specialized models, it allows seamless “plug-and-play” integration of new or improved expert systems as they emerge. Moreover, it supports the incorporation of external knowledge ($k$), such as textual descriptions of chemical rules. Consequently, this design ensures the framework’s long-term adaptability and obviates the need for costly retraining of the LLM when integrating new components.
Collaborative Reasoning Engine
@@
Building upon the chemical decision space, the Collaborative Reasoning Engine acts as the framework’s cognitive hub, bridging the shallow, heuristic-driven inference of specialized models with the deep, logic-based reasoning of the LLMs. This collaborative process redefines the role of the LLM, transforming it from a mere predictor or tool-caller into an active reasoning agent that navigates the full retrosynthetic path. Concurrently, specialized models also transition from being passive data providers to active executors of shallow reasoning, tasked with constructing the actionable space. This profound cognitive collaboration guides the LLM to shift from “directly predicting reactants” to “step-by-step logical deduction” upon the constructed chemical reasoning space.
The collaborative mechanism is implemented through two key steps: (1) Reasoning Context Construction. This initial step builds a structured reasoning environment for the LLM. Based on the chemical decision space, we organize the candidate set for each sub-task following a logical hierarchy. In this hierarchy, the global decision (e.g., selecting the overall reaction type) is addressed first because it critically guides the subsequent search for local specifics (e.g., identifying the precise reaction center). This hierarchical organization yields $n$ interconnected, difficulty-increasing candidate sets $\mathcal{T}$. These candidate sets, combined with any external domain knowledge $k$, collectively form the LLM’s reasoning context. The resulting prompt is thus designed to be a structured yet open-ended environment, guiding the LLM while encouraging it to leverage both the curated candidates and its own intrinsic knowledge. (2) LLM-driven Critical-Generative Reasoning. Within this tailored context, the LLM operates as the active chemical reasoner. It leverages the provided information to autonomously execute a step-by-step reasoning chain, moving sequentially through sub-tasks to construct a logically coherent retrosynthetic pathway.
@@
Specifically, the LLM’s autonomous reasoning is a dynamic interplay of two core capabilities: critical analysis and generative decision-making. At each reasoning step, the LLM first rigorously evaluates the candidate set for the current sub-task within the provided context. This critical analysis, which ensures a comprehensive exploration of the problem space, directly informs the subsequent decision-making process, which manifests in one of two actions: (1) Selection: The LLM identifies and selects the most plausible candidate from the provided options. (2) Generation: If its analysis concludes that no provided candidate is satisfactory, the LLM leverages its internal knowledge and reasoning context to generate a novel, self-consistent solution. This adaptive reasoning strategy, by mirroring human expert intuition, is therefore not only more efficient and context-aware but also fundamental to enhancing both the accuracy and explainability of the final prediction.
@@
T5Chem (Lu and Zhang 2022) 43.64 0.972 8.42 0.982 0.956 0.974 0.925
ChemFormer (Irwin et al. 2022) 27.30 0.769 14.77 0.952 0.782 0.690 0.647
nach0 (Livne et al. 2024) 30.52 0.976 11.91 0.990 0.947 0.977 0.905
ChemLLM (Zhang et al. 2024) 12.16 0.868 10.42 0.977 0.849 0.846 0.772
ChemDFM (Zhao et al. 2024) 6.69 0.743 16.62 0.874 0.714 0.807 0.617
ChemCrow (+nach0) (Bran et al. 2023) 30.52 0.976 11.91 0.990 0.947 0.977 0.905
@@
GPT-3.5 31.30 0.967 7.87 0.991 0.949 0.964 0.910
GPT-4o 31.21 0.968 7.80 0.991 0.950 0.965 0.912
InterLM-2-Chat-7B 23.12 0.954 11.17 0.893 0.944 0.951 0.907
Gemini-2.5-preview-05-20 37.11 0.984 6.75 0.959 0.971 0.988 0.942
Qwen2.5-7B-Instruct 36.09 0.970 7.77 0.992 0.958 0.972 0.920
Retro-Expert 66.23 0.995 4.41 0.997 0.986 0.994 0.972
Table 1: Comparison of LLM-based models on the retrosynthesis task. * the model is provided with Top-4 reactant candidates and additional domain knowledge anchors (i.e., reaction types, reaction center) from specialized models. Results for other models are reproduced following their original evaluation methods.
Knowledge-Guided Policy Optimization
To overcome the limitations of SFT, which primarily trains models to memorize common reaction patterns from datasets, we introduce a reinforcement learning framework to optimize the LLM’s reasoning policy within the constructed chemical decision space. At the heart of this framework lies our proposed Knowledge-Guided Policy Optimization (KGPO) mechanism. The core objective of KGPO is to incentivize the LLM to learn a chemically-grounded reasoning policy that prioritizes the logical validity of the entire pathway, not merely the correctness of the final prediction. The objective function for this optimization is formally defined as follows:

### v2

@@ Abstract
Retrosynthesis prediction aims to infer the reactant molecule based on a given product molecule, which is a fundamental task in chemical synthesis. However, existing models rely on static pattern-matching paradigm, which limits their ability to perform effective logic decision-making, leading to black-box decision-making. Building on this, we propose Retro-Expert, an interpretable retrosynthesis framework that performs collaborative reasoning by combining the complementary reasoning strengths of Large Language Models and specialized models via reinforcement learning. It outputs natural language explanations grounded in chemical logic through three components: (1) specialized models analyze the product to construct high-quality chemical decision space, (2) LLM-driven critical reasoning to generate predictions and corresponding interpretable reasoning path, and (3) reinforcement learning optimizing interpretable decision policy. Experiments show that Retro-Expert not only surpasses both LLM-based and specialized models across different metrics but also provides expert-aligned explanations that bridge the gap between AI predictions and actionable chemical insights.
1 Introduction
@@
Retrosynthesis prediction aims to deduce potential reactants and reaction pathways for synthesizing a target product molecule based on its structural characteristics [30, 25, 32], holding significant application value in drug discovery and molecular design [14, 38, 35]. However, existing models primarily adopt a pattern-matching paradigm, which learn mappings between product SMILES and reactant SMILES from datasets, framing the task as either classification or auto-regressive sequence generation [34, 4, 46]. Under this paradigm, traditional specialized models capture intricate structural features to learn improved mappings, yet their black-box prediction mechanism inherently limits their ability to provide chemically-grounded rationales for their outputs. In contrast, large language models (LLMs) rely on supervised fine-tuning (SFT) to enhance retrosynthetic capabilities, but their performance depends heavily on memorized reaction patterns and typically outputs only reactants or corresponding options [44, 45]. This reliance limits their capacity to understand underlying chemical principles and perform logical reasoning. This paradigm faces three key limitations: (1) The model only generates reactant SMILES strings, with no transparency in its internal reasoning process. (2) The predictions lack natural language explanations grounded in chemical logic, critically hindering real-world adoption. (3) The pure pattern-matching mechanism significantly constrains the model’s capacity for effective logical decision-making. These limitations lead to a lack of reliable basis for the predictions in chemical principles, which severely undermines chemists’ trust in practical applications.
Figure 1: Comparison between the conventional RL pipeline and our methods. Conventional methods require SFT as a necessary step for effective RL, whereas our methods directly optimizes the LLM by leveraging the decision space.
@@
Given the mentioned challenges, a reasoning-driven paradigm aligns more closely with the intrinsic nature of retrosynthetic analysis than simplistic pattern matching. This approach mirrors the strategy employed by chemical experts, who conceptualize retrosynthesis as an iterative and logical reasoning process. Notably, recent breakthroughs in other domains [36, 39, 5] have demonstrated the potential of LLMs to address complex specialized problems through specialized-knowledge-based reasoning, enhanced by reinforcement learning. These findings motivate us to explore how to leverage LLMs’ emergent reasoning capabilities to enhance the interpretability of retrosynthesis prediction.
@@
We focus on chemical knowledge-based retrosynthetic reasoning by LLMs to generate reactants along with an explainable reasoning process, ensuring interpretable and transparent retrosynthesis prediction. To induce reasoning, the selection of an appropriate learning paradigm is crucial. Supervised Fine-Tuning (SFT) inclines models towards replicating prevalent reaction patterns observed in training data, rather than engaging in reasoning grounded in underlying chemical principles. In contrast, Reinforcement Learning (RL) approaches, which optimize the models’ behavioral policy through reward feedback, offer a theoretically viable path to incentivize the complex and logical reasoning abilities for chemical tasks.
@@
However, directly applying RL to incentivize models for retrosynthetic reasoning faces two critical challenges: (1) Domain Knowledge Disparity. Retrosynthesis demands not only logical reasoning but also mastery of specialized chemical knowledge. Pre-trained LLMs fail to adequately internalize and apply specific chemical principles when reasoning solely based on molecular SMILES. (2) Granularity Mismatch in Specialized Capabilities. The retrosynthetic reasoning process comprises multiple subtasks with distinct cognitive requirements, including reaction type classification and reaction center localization [9, 37, 40]. While general LLMs demonstrate profound logical reasoning abilities, they inherently lack precise perception and matching capabilities for fundamental chemical reaction patterns - a critical requirement for certain subtasks. In contrast, dedicated specialized models, each individually optimized for a specific subtask, have achieved strong performance [38, 4, 30]. To bridge this gap, we propose a novel paradigm that synergistically combines: (1) the knowledge grounding capabilities of specialized models with (2) the deep logical reasoning of LLMs. Specifically, specialized models construct high-quality, multi-dimensional chemical decision spaces, upon which LLMs perform logical reasoning to make the prediction process interpretable.
@@
Building upon these insights, we present Explainable and Cooperative retrosynthesis framework, Retro-Expert, the first explainable retrosynthesis framework that integrates natural language-based expert reasoning with model-agnostic compatibility (in Figure 1). The framework contains three synergistic core components: (1) Chemical Decision Space Construction. Leveraging specialized models (e.g., reaction type classifiers, reaction center localization and reactant generators) to construct high-quality chemical decision space, which provides “knowledge anchors” for LLM’s subsequent deep reasoning. (2) Collaborative Reasoning Engine. Leveraging the effective collaboration with specialized models, the LLM perform integrated critic analysis and in-depth reasoning upon the chemical decision space. It produces the final reactant prediction along with interpretable, natural language-based reasoning process. (3) Knowledge-Guided Policy Optimization. Retro-Expert optimizes the LLM’s reasoning strategy via rule-based reinforcement learning. A multi-stage reward mechanism is designed during training to guide the model toward learning an optimal and trustworthy reasoning path. Notably, Retro-Expert allows seamless integration of arbitrary specialized models during inference, enabling flexible expansion and adaptation without requiring retraining.
@@
2. We propose a collaborative retrosynthesis reasoning framework built upon the synergy between the LLM and specialized models. It improves prediction accuracy while generates human-understandable analysis.
@@
2 Related Work
Retrosynthesis Prediction. Existing retrosynthesis methods are broadly categorized into two modeling paradigms: specialized machine learning (ML) models and large language model (LLM)-based methods. ML-based approaches model chemical molecules from different perspectives (e.g., SMILES strings, molecular graphs) to learn the mapping from product to reactant molecules [23, 48, 4, 30]. Although these methods demonstrate significant potential through precise molecular mappings, their black-box decision-making mechanisms inherently restrict the model’s interpretability.
LLM-based approaches aim to enhance the comprehension and memorization capabilities of LLMs for chemical tasks. Some efforts tailor LLMs to retrosynthesis via SFT [44, 45, 42]. Limited by the memorization-centric nature of SFT, these models tend to replicating memorized reaction patterns and only produce predicted reactants or corresponding options. Other work integrates the LLM with external tools to improve prediction accuracy [1, 2, 18]. However, they are limited to directly adopting or scoring based on predefined metrics of specialized model predictions. Such strategies lack engagement in the predictive process and thus fail to provide interpretable reasoning pathways that reveal underlying reaction mechanisms.
Large Language Model Reasoning. In recent years, the profound reasoning capabilities of large language models (LLMs) have been developed to solve specialized scientific problems [31, 33, 22, 21]. While employing SFT [49, 15], the model primarily learns to replicate common reasoning patterns from curated datasets, restricting its effectiveness on more complex tasks. In contrast, reinforcement learning (RL)-based reasoning models recently achieve significant progress [11, 50, 8]. In the chemistry field, some studies have assessed the performance of LLMs on chemical reasoning tasks [12, 7]. These models predominantly focus on enhancing performance through SFT paradigm. However, there are few studies that incorporate expert-level reasoning strategies to tackle complex chemical reasoning tasks.
@@ 2 Related Work
3 Methodology
3.1 Overview
Retro-Expert is to enhance prediction accuracy while generating human-understandable reasoning process. As illustrated in Figure 2, Retro-Expert is built upon three core modules. The process starts with Chemical Decision Space Construction, where specialized models analyze the target product and construct an anchored chemical decision space. This foundation activates the Collaborative Reasoning Engine, where the LLM performs deep reasoning by critically analyzing candidates, selecting an optimal path, and articulating its multi-step logic into a complete natural language chain. Finally, Knowledge-Guided Decision Policy Optimization employs RL to refine the LLM’s decision-making strategy, using a multi-stage reward mechanism to guide it towards an optimal and trustworthy reasoning path.
3.2 Problem Formulation
The objective of the retrosynthesis task $T_{\text{retro}}$ is to predict the set of reactants $\{M_{r}^{i}\}_{i=1}^{C}(C\geq 1)$ corresponding to a target product $M_{p}$. According to chemical expert knowledge, the overall task can be decomposed into $s$ logically connected subtasks $T_{\text{retro}}=\{T_{0},T_{1},\dots,T_{s}\}$. Given an environment $\mathcal{E}$ containing $S(S\geq s)$ specialized models $\mathcal{M}=\{m_{0},m_{1},\dots,m_{S}\}$, each model $m_{i}$ performs experience-based domain-specific reasoning to generate Top-$N$ candidate predictions $P_{i}$ for its corresponding subtask:
$P_{i}=\{P_{i}^{n}\}_{n=1}^{N},\quad P_{i}^{n}\sim p(m_{i}|M_{p};\theta),$ (1)
where $\theta$ denotes the parameters of model $m_{i}$, and $P_{i}^{n}$ represents the $n$-th candidate result. The chemical decision space $\mathcal{T}$ is defined as the Cartesian product of candidate results from all subtasks:
$\mathcal{T}=(P_{0},P_{1},\dots,P_{s}),$ (2)
with a space size of $K^{s}$. Upon this space, an LLM $M_{\text{LLM}}$ interacts with $\mathcal{E}$, critically analyzes the Top-$N$ candidates from models in $\mathcal{M}$, and infers the correct answer $P_{i}^{\prime}$ for the subtask. This generates a reasoning path $\mathcal{T}_{\text{LLM}}=(P_{0}^{\prime},P_{1}^{\prime},\dots,P_{s}^{\prime})$ alongside a natural-language explanation $R$. Retro-Expert has two core objectives: generating the correct set of reactants $\hat{a}=\{M_{r}^{i}\}_{i=1}^{C}$, and identifying the optimal path $T^{*}$ with the highest reward within $\mathcal{T}$:
@@ 3.2 Problem Formulation
3.3 Chemical Decision Space Construction
Due to the lack of domain-specific knowledge and experience, general LLMs face challenges in retrosynthesis when relying solely on molecular SMILES. Existing methods rely on SFT on large structured chain-of-thought (CoT) data for effective RL. Meanwhile, directly applying RL often fails to efficiently explore feasible reasoning paths under limited sampling, making it difficult for training to progress.
@@ 3.3 Chemical Decision Space Construction
To overcome these challenges, we propose the chemical decision space, which provides LLMs with chemically meaningful starting points for reasoning. This design enables the model to explore feasible reaction pathways more efficiently while maintaining flexibility in its reasoning process. Specifically, for a given target product $M_{p}$, we first decompose the retrosynthesis task into a set of distinct chemical sub-tasks such as reaction type prediction and reaction center localization. We then invoke a suite of specialized models optimized for high recall, each dedicated to a sub-task, to generate a corresponding set of plausible candidates. These candidate sets are then integrated, with each subtask’s knowledge (e.g., reaction type, reaction center) forming a multi-dimensional decision space. This space provides well-founded “knowledge anchors” that ground the LLM’s subsequent reasoning, enabling knowledge-aware strategic planning and efficient exploration.
3.4 Collaborative Reasoning Engine
@@
Building upon the chemical decision space, the Collaborative Reasoning Engine acts as the framework’s cognitive hub, bridging specialized models with the LLM. This collaborative process shifts the LLM’s role from a mere predictor or tool-caller to an active reasoning agent that navigates the full retrosynthetic path. This cognitive collaboration guides the LLM to shift from “directly predicting reactants” to “logical deduction” upon chemical decision space.
The collaborative mechanism is implemented through two key steps: (1) Reasoning Context Construction. Based on the chemical decision space, we combine the candidate sets for each sub-task with external knowledge $L$ to form part of the LLM’s reasoning input. (2) LLM-driven Critical-Generative Reasoning. Within the context, the LLM acts as the active chemical reasoner. In contrast to sequential subtask execution, the LLM autonomously performs logical reasoning by selectively utilizing provided candidate sets, constructing a logically coherent retrosynthetic pathway.
@@
Specifically, the LLM’s autonomous reasoning is a dynamic interplay of two core capabilities: critical analysis and generative decision-making. At each reasoning step, the LLM first rigorously evaluates one candidate set within the provided context. This critical analysis, which ensures a comprehensive exploration of the problem space, directly informs the subsequent decision-making process, which manifests in one of two actions: (1) Selection: The LLM identifies and selects the most plausible candidate from the provided options. (2) Generation: If its analysis concludes that no provided candidate is satisfactory, the LLM leverages its internal knowledge and reasoning context to generate a novel, self-consistent solution. This adaptive reasoning strategy, by mirroring human expert intuition, is therefore not only more efficient and context-aware but also fundamental to enhancing both the accuracy and explainability of the final prediction.
Table 1: Comparison with LLM-based models on USPTO-50K [24]. Space Setting indicates whether the model is provided with Top-4 reactant candidates and additional domain knowledge anchors (i.e., reaction types, reaction center) from specialized models. General LLMs excel at logical reasoning but lack chemical knowledge. When equipped with a chemical decision space, their performance improves substantially. Results of all LLMs (with/without space) are reported in the Appendix.
@@
T5Chem [20] 43.64 0.972 8.42 0.982 0.956 0.974 0.925
ChemFormer [16] 27.30 0.769 14.77 0.952 0.782 0.690 0.647
nach0 [19] 30.52 0.976 11.91 0.990 0.947 0.977 0.905
ChemLLM [44] 12.83 0.869 10.34 0.979 0.851 0.849 0.773
ChemDFM [45] 6.69 0.743 16.62 0.874 0.714 0.807 0.617
ChemCrow (+nach0) [1] 30.52 0.976 11.91 0.990 0.947 0.977 0.905
@@
InterLM-2-Chat-7B 24.31 0.957 11.03 0.895 0.945 0.952 0.909
GPT-3.5 32.93 0.969 7.76 0.991 0.952 0.968 0.913
GPT-4o 32.85 0.970 7.72 0.990 0.953 0.970 0.914
Gemini-2.5-preview-05-20 39.02 0.986 6.63 0.961 0.973 0.989 0.944
Qwen2.5-7B-Instruct [41] 37.55 0.971 7.64 0.992 0.961 0.973 0.922
Ours 70.28 0.998 2.27 0.997 0.990 0.996 0.989
3.5 Knowledge-Guided Policy Optimization
To overcome the limitations of SFT, we introduce a reinforcement learning framework to optimize the LLM’s reasoning policy within the constructed chemical decision space. The objective function is formally defined as follows: